# Process Reinforcement through Implicit Rewards

### v1

@@ 5.2 Reference Model Choice is Flexible
5.3 Single-Forward v.s. Double-Forward
@@
5.4 PRIME with Other RL Algorithms
As we stated before, PRIME is equally applicable to other RL algorithms beyond RLOO. In this section, we implement PRIME with REINFORCE (Williams, 1992), GRPO (Shao et al., 2024), and PPO (Schulman et al., 2017). Similarly to RLOO, we only modify the advantage estimation functions and leave the clip surrogate loss unchanged.
For REINFORCE, we have
From Figure 11 and Table 3, We show that PRIME boosts these algorithms on both efficiency and performance as it does with RLOO. PRIME contributes consistently regardless of the policy update method, making it a generic algorithm. It indicates that PRIME is a general plug-in for almost any RL algorithm for LLM., which largely extends the use cases of PRIME.
Moreover, the PPO variant of PRIME provides no performance gain, demonstrating that the additional computation cost from the critic model is redundant. This makes it possible to compensate for the expense of the process reward model by using REINFORCE-like algorithms with simpler advantage estimators. Finally, we choose the best-performing RLOO as the advantage estimator in our algorithm.
@@ 5.4 PRIME with Other RL Algorithms
Figure 10: PRIME also benefits REINFORCE, GRPO, and PPO, achieving similar improvement as RLOO.
Figure 11: Comparison of value models and reward models. We show that value models, either the original PPO one or Implicit PRM, is substaintially worse than reward models.
5.5 Value or Reward, How to Use the Implicit PRM?
(1) REINFORCE: $A_{t}=r_{o}(\mathbf{y})$.
(2) On top of (1), using a linear-head value model $V$ to calculate the baseline: $A_{t}=r_{o}(\mathbf{y})-V(\mathbf{y}_{<t})$. This is the original PPO in Figure 11 as we set $\gamma=1$ and $\lambda=1$.
(3) On top of (1), using values from the Implicit PRM to serve as the baseline: $A_{t}=r_{o}(\mathbf{y})-v_{\phi}(\mathbf{y}_{<t})$. This is equivalent to PPO with its value model being replaced by values from the Implicit PRM when $\gamma=1$ and $\lambda=1$.
(4) On top of (1), using process rewards from the Implicit PRM to calculate the return: $A_{t}=r_{o}(\mathbf{y})+\sum_{s=t}^{T}r_{\phi}(y_{s})$. This is the REINFORCE w/ PRIME in Figure 11.
Figure 11 reports the results. Comparing PPO and REINFORCE, we find that an additional value model does not benefit policy performance. Notably, using rewards from the Implicit PRM to calculate returns, which is the default setting in PRIME, greatly outperforms all three baselines, regardless of where the values come from. This indicates that PRMs work better than value models in RL for LLMs.
5.6 “Zero” Experiments
@@ 5.5 Value or Reward, How to Use the Implicit PRM?
6 Related Work
RL for LLM Reasoning. In the context of LLMs, reinforcement learning has been widely used for aligning human preferences (Christiano et al., 2017; Ouyang et al., 2022; Cui et al., 2024), but the open-source community mostly adopt the data-driven imitation learning methods (Yuan et al., 2024a; Yue et al., 2024; Wei et al., 2024; Liu et al., 2024) to enhance the reasoning capabities of LLMs. Over the past few months, the paradigm gradually shifted. OpenAI o1 (Jaech et al., 2024) first showed the tremendous potential of large-sacle RL for reasoning LLMs, and recent works have verified the scaling effect of the simple RL recipe with merely outcome rewards (DeepSeek-AI et al., 2025; Team et al., 2025). Meanwhile, the role of dense rewards in RL remains underexplored, which is the main focus of PRIME.
Implicit Rewards. Implicit rewards are broadly adopted in LLM alignment (Rafailov et al., 2023; Chen et al., 2024b; Azar et al., 2024; Ethayarajh et al., 2024; Rosset et al., 2024; Chen et al., 2024a).  Rafailov et al. (2024) first showed that optimizing DPO objective learns a Q function implicitly. Zhou et al. (2024) utilized implicit rewards in PPO, and showed that dense implicit rewards are better than sparse ones.  Yuan et al. (2024b) further extended the conclusion to any loss funtion optimizing Eq. 3.
7 Conclusion
As the fuel of LLMs, data, will be depleted in the near future, we are entering a new era of search and exploration, which is exemplified by reinforcement learning (Sutton, 2019). This work develops PRIME, which produces and leverages dense rewards in online RL for LLM reasoning. Throughout the experiments, we validate that PRIME (1) greatly benefits sample efficiency and policy performance, (2) is easy to use with minimum cost, and (3) is a general method that works with broad RL algorithms together.
Appendix A SFT Data & Training Details
@@
We first perform supervised finetuning on the base model to get a starter model for RL.
@@
Appendix B RL Data Preprocessing
B.1 RL Data Collection and Preprocessing
@@
B.2 Data Filtering and Question-Type Classification
@@
B.3 Converting to Direct Question-Answer Format
@@
B.4 Problem and Solution Validation

### v2

@@ B.1 Reference Model Choice is Flexible
B.2 Single-Forward v.s. Double-Forward
@@
Model Eurus-2-7B-PRIME Qwen2.5-Math-7B-Instruct
Base Model Qwen2.5-Math-7B Qwen2.5-Math-7B
SFT Data 230K (open-source) 2.5M (open-source & in-house)
RM Data 0 618K (in-house)
RM Eurus-2-7B-SFT Qwen2.5-Math-RM (72B)
RL Data 150K queries $\times$ 4 samples 66K queries $\times$ 32 samples
B.3 Results of Different RL Algorithms
@@
We ablate PRIME and different RL algorithms with their variants and find that the PRIME algorithm achieves the best performance for several reasons.
For REINFORCE, we have
As shown in Table 4, PRIME contributes consistently regardless of the policy update method, making it a generic algorithm.
Moreover, the PPO variant of PRIME provides no performance gain, demonstrating that the additional computation cost from the critic model is redundant. This makes it possible to compensate for the expense of the process reward model by using REINFORCE-like algorithms with simpler advantage estimators.
Finally, we choose the best-performing RLOO as the advantage estimator in our algorithm.
B.4 “Zero” Experiments
@@
B.5 Effect of Reward Model Size
In the main experiments, we set the reward model to be of the same size as the policy model by default. To further investigate the influence of reward model capacity, we conduct comparative experiments by fixing the policy model as Qwen2.5-7B-Base and varying the reward model among Qwen2.5-3B-Base, Qwen2.5-7B-Base, and Qwen2.5-14B-Base, other settings are aligned with the main experiment. The results are summarized in Table 5. Overall, the results suggest that reward model size has limited influence: the 7B reward model achieves the best average performance, while larger (14B) or smaller (3B) reward models do not yield clear advantages.
Table 5: Performance comparison of different reward models (Qwen2.5-7B-Base as policy model).
Reward Model AIME 24 AIME 25 AMC MATH Minerva OlympiadBench Average
Qwen2.5-3B 10.7 4.8 44.0 73.2 26.1 33.0 32.0
Qwen2.5-7B 13.2 6.4 42.9 73.4 26.5 33.1 32.6
Qwen2.5-14B 10.8 4.8 44.1 73.2 25.4 32.7 31.8
B.6 Comparison with VinePPO
VinePPO (Kazemnejad et al., 2024) uses average return across trajectories to estimate the value within policy gradient update, instead of using value model in algorithm like PPO. We adopt the same setting as VinePPO, using RhoMath 1.1B (Lin et al., 2024) as the base model and MATH dataset for training. We reproduced PRIME and VinePPO with 8 A800-80G for 96 steps with the same hyperparameters. It reveals that
(1) PRIME is 11x more efficient than VinePPO. As shown in Figure 16, VinePPO takes 13.94 hours for training, while PRIME only needs 1.22 hours.
(2) PRIME also consistently outperforms VinePPO on the validation set. As shown in the Table 6, the validation accuracy of PRIME is higher than VinePPO at each validation step.
Table 6: Comparison between PRIME and VinePPO.
Steps 16 32 48 64 80 96
VinePPO Val Acc (%) 15.7 16.3 17.2 17.6 17.7 18.4
VinePPO Clock Time (Hours) 2.23 4.57 7.23 9.86 11.96 13.94
PRIME Val Acc (%) 16.4 16.8 17.5 18.1 18.7 18.8
PRIME Clock Time (Hours) 0.22 0.41 0.60 0.80 1.01 1.22
Figure 15: Validation accuracy curves of PRIME and VinePPO.
Figure 16: Training reward curves of PRIME and DeepScaleR.
Figure 17: The effect of dense reward. We compared PRIME and RLOO with outcome verifier (OV). The figure depicts training reward curves across wall clock, revealing better sample efficiency of PRIME.
B.7 Comarison with DeepScaleR
DeepScaleR (Luo et al., 2025) introduces a three-stage training pipeline, with maximum allowed response length iteratively increased, from 8k (Stage 1) to 16k (Stage 2), finally to 24k (Stage 3). Through this iterative context lengthening, DeepScaleR achieve continuous performance gain. We run PRIME under the setting of DeepScaleR (base model, data, hyperparameters strictly follow the official settings). Due to resource limits, we only conduct the first (8k) stage training for 330 steps, yet PRIME achieves impressive results.
(1) Model Performance. Figure 16 presents the training accuracy curves of PRIME and DeepScaleR [^1] Note that the provided training logs are broken according to this issue, so we estimated its training curve from the figure in the blog.. PRIME achieves comparable training accuracy within 330 steps, which is only 1/3 the steps of DeepScaleR stage 1 (1040 steps). On testsets, as shown in Table 7, PRIME consistently improves the performance of DeepSeek-R1-Distill-Qwen-1.5B by 3.1 points. This validates the effectiveness of PRIME on highly capable base models.
(2) Efficiency. PRIME consumed 446.7 A800 GPU hours for this experiment. In contrast, DeepScaleR consumed 3800 A100 GPU hours in total, and the first stage roughly required $\sim$600 GPU hours. This means PRIME is also 25% faster than DeepScaleR. Note that the advantage could be higher considering hardware differences (A800/A100).
(3) Computation Overhead. Moreover, due to the long response length of the distill model, the overhead would be attributed more to the generation phase, narrowing down the extra time PRIME brings to about 18%. This means that PRIME would be more suitable for long reasoning models.
Table 7: Comparison between PRIME and DeepScaleR.
Model Step GPU Hour AIME 2024 MATH-500 AMC MinervaMath OlympiadBench Avg.
DeepScaleR-1.5B-Preview 1750 3800 43.1 87.8 73.6 30.2 50.0 57.0
DeepScaleR-1.5B-Stage1 1040 $\sim$ 600 33.9 - - - - -
DeepSeek-R1-Distill-Qwen-1.5B - - 28.8 82.8 62.9 26.5 43.3 48.9
PRIME-DeepScaleR-1.5B-Stage1 330 446.7 32.1 85.1 68.1 30.1 44.6 52.0
Appendix C Discussion on Implicit Process Reward
C.1 Formulation Validity
As shown in Yuan et al. (2024b), implicit process reward is a parameterization of reward modeling. With such parameterization, the expectation of cumulative reward starting from the $y_{t}$ (i.e. q-value), $q_{\phi}^{t}(\mathbf{y}_{<t},y_{t})=\beta\log\mathbb{E}_{\pi_{\text{ref}}(\mathbf{y}|\mathbf{y}_{\leq t})}e^{\frac{1}{\beta}r_{\phi}(\mathbf{y})}$ would have closed-formed solution $q_{\phi}^{t}(\mathbf{y}_{<t},y_{t})=\sum_{i=1}^{t}\beta\log\frac{\pi_{\phi}(y_{i}|\mathbf{y}_{<i})}{\pi_{\text{ref}}(y_{i}|\mathbf{y}_{<i})}$. Therefore, despite the similarities between implicit process reward formulation to those of DPO (Rafailov et al., 2023; 2024), it is not derived from the optimal policy of entropy-regularized RL (Ziebart et al., 2008; Haarnoja et al., 2017).
C.2 Loss Function
Since $\pi_{\phi}$ and $\pi_{\text{ref}}$ are language models which are inherently self-normalized, using cross-entropy loss brings about a minor issue that the minimum loss 0 cannot be reached. Therefore, the optimal solution would satisfy $\beta\log\frac{\pi_{\phi}^{*}(\mathbf{y})}{\pi_{\text{ref}}(\mathbf{y})}=r_{o}+c$ rather than $\beta\log\frac{\pi_{\phi}^{*}(\mathbf{y})}{\pi_{\text{ref}}(\mathbf{y})}=r_{o}$. This discrepancy would not affect RL, because PRIME uses relative reward, $r_{\phi}\left(\mathbf{y}^{i}\right)-\frac{1}{K-1}\sum_{j\neq i}r_{\phi}\left(\mathbf{y}^{j}\right)$, rather than the original reward from PRM. This means that even if the bias term is included, it would be canceled out in calculation since the bias term is only related with prompt $x$.
To solve this issue, we can simply eliminate this value term by using the DPO loss. We have a pilot experiment comparing DPO and CE loss, as shown in Table 8 and Figure 18. DPO and CE achieve similar results, and we chose CE for memory efficiency.
Table 8: Test accuracy of updating PRM with CE or DPO loss after training.
Method Step AIME 2024 AMC MATH-500 MinervaMath OlympiadBench Avg.
PRIME w. DPO loss 96 7.7 39.3 66.2 17.3 31.3 32.4
PRIME w. CE loss 96 7.9 40.2 66.0 16.9 30.7 32.3
(a) Outcome training rewards.
(b) Math test accuracy across different gradient steps.
Figure 18: Outcome rewards and test accuracy of updating PRM with DPO or CE loss during training.
C.3 Reward Shaping
Another view of PRIME is potential-based reward shaping (Ng et al., 1999). If we view the q-value as the potential function, the process reward exactly satisfies the definition of shaping reward ($\gamma=1$). PBRS does not affect the optimal policy, but does speed up learning, which is aligned with our results.
Appendix D SFT Data & Training Details
@@
We first performed supervised finetuning for RL preparation. We focus on mathematical and coding problems in this paper. For models, we start with Qwen2.5-Math-7B-Base (Yang et al., 2024b) for its great mathematical capabilities.
@@
Appendix E RL Data Preprocessing
E.1 RL Data Collection and Preprocessing
@@
E.2 Data Filtering and Question-Type Classification
@@
E.3 Converting to Direct Question-Answer Format
@@
E.4 Problem and Solution Validation
@@ E.5 PRM Data
Table 12: Avg@16 results with temperature=0.3 of PRIME and RLOO w/ outcome verifier (OV).
Method Step AIME 2024 AMC
Eurus-2-7B-SFT 0 4.4 21.4
RLOO w/ OV Only 240 15.4 43.8
Eurus-2-7B-PRIME 240 17.3 49.2
Eurus-2-7B-PRIME 592 24.2 54.5